film of refractive index 1.50 and


thickness 0.25 mm is introduced in the


path of one of the interfering beams of a


Michelson's interferometer. It causes


600 dark fringes to sweep across the


field of view. Calculate the wavelength


of light used.

1
Expert's answer
2022-03-15T16:49:59-0400

For wavelength in this case


"\lambda = \frac{2d(n-1)}{z}"

Calculating

"\lambda=\frac{2\cdot0.25\cdot(1.5-1)}{600}=416" nm